you are out jogging in the city. you start out by running 8 miles north and 3 miles south. after you run 3.5 mles east you come across a street that will take you straight back home. you decide that is enough jogging for the day and decde to walk home on the straight line street. how far do you have to walk?
Answer by CubeyThePenguin(3113)

displacement: (8 - 3) = + 5 miles in the vertical direction
3.5 miles in the horizontal direction

miles walked = sqrt(5^2 + 3.5^2) = 6.1 miles